What is the total surface area of a hockey puck with diameter 7.2cm and height
2.3cm?

Answer by nerdybill(7384)

What is the total surface area of a hockey puck with diameter 7.2cm and height
2.3cm?
.
radius = 7.2/2 = 3.6
area of top and bottom:
2(pi)r^2 = 2(3.14)(3.6)^2 = 81.38 square cm
.
area of side:
(pi)7.2 = (3.14)7.2 = 22.61 square cm
.
total surface area: 104 square cm